OpenStudy (anonymous):

Compute the volume change of a solid copper cube, 40mm on each edge, whaen subjected to a pressure of 20Mpa. The bulk modulus of copper is 125Gpa.

OpenStudy (anonymous):

B=PdV/V implies dV=BV/P Simply substitute the values We get dV= \[\frac{ 125*10^{9}*(40*10^{-3})^{3} }{20*10^{6} }\]